Economics class

How much will Kingston technologies have to pay each year in 7 equal payments, starting 2 years from now, to repay a $900,000 loan. The interest rate is 9 % per year?


Kingston Technologies will have to pay $______ each year to repay the loan 
Answer #1

Interest. Rate per year is = 9%

Total installment made per year is 7 which is = 9/7=1.28 % per installments

Total principle amount is $900000 which is divide Into 14 equal installments. $900000/14=64285.7143 per installment without interest

Interest amount will be 1.28*9000=$11520

Total amount pay per installment is $64285.7143+$11520=$75805.7143

Therefore amount per year is $75805.7143*7= $530640

hence, $530640 amount required to repay loan per year


answered by: doge